Calculate how many exterior studs will be needed; How much Exterior wall plate stock will be needed (add 5% for waste); lineal footage of door & window headers (based off of the concepts learned in class, add 5% for waste); lineal footage of rough window sills (add 5% for waste); lineal footage of header caps ( add 5% for waste); how many sheets of sheathing will be needed (add 5% for waste); calculate one brace per 20' and two for each corner. The walls will be constructed of 2" x 6" material, the header material will be 4" x 10" & the plywood to be used will be 7/16" O.S.B.

House #1: the dimensions of the structure are 48'- 0" wide X 24'- 0" in depth; the plate height for the walls is 8'-1"; there will be 3-4'-0" windows; 1-3'-0" X 6'-8" door; and the floor joists are set down in the foundation wall.

Number of studs:_____________________________________

Plate stock:__________________________________________

Header stock:_________________________________________

Rough sill stock:_______________________________________

Header cap stock:_____________________________________

Sheets of sheathing:___________________________________

Number of wall braces:_________________________________
Answer #1

Number of Studs

Wall length* 0.75(center stud with 16" default parameter)

= 48'*24'

= 1152

= 1152*0.75= 864

Total number of exterior walls stud= 864

Plate stock

The total length of wall*3 times

= 48'*3= 144

Add: 5% waste(144*5%) = 7.2

Plate stock= 144+7.2=151.2

Header Stock

Window opening

= 3'4"+7"

= 3'11"

Doors

6'8"+7"

= 7'4"

Add: 5% waste = 3'11*7'4

= 24.1647

Rough sill stock

= 5%(3*5/100)

= 0.0125"

Header Cup stock

= 4*10

= 40

Add waste 5%= 40*5%= 8

= 40+8= 48 inches

Sheets of sheathing

Area of wall= 48'*24'

= 1152

Area of the window opening

= 3'*4'

= 12

Area of the Door opening

= 6'8"*1'3"

= 106.08

The total area of wall-total area of the opening

= 1152-106.08

= 1045.2

Number of sheathing sheets= 1045.2/32

= 32.66

= 33 sheets(Roundup)

Number of wall braces

= total wall area/required length per brace

= 1152/20

= 57.6

= 56 wall braces